OpenStudy (anonymous):

what is an equation for the line that passes through points (-1 -4) and (1, 4)? Write the equation in slope-intercept form

OpenStudy (zephyr141):

\[y-y_{1}=m(x-x_{1})\] just plug it in.

OpenStudy (agent0smith):

First find the slope \[m = \frac{ y_2-y_1 }{ x_2 - x_1 }\] then plug into the formula zephyr gave

OpenStudy (anonymous):

idk what goes were...

OpenStudy (agent0smith):

you were given two points, (x1, y1) and (x2, y2) which are (-1 -4) and (1, 4) so plug those into the slope formula

OpenStudy (anonymous):

4^2--(4) / 1-(-1)

OpenStudy (agent0smith):

correct, but you have a ^2 there which should not be